{F} Calculate the approximate value of 10 to four decimal places by taking the first four


terms of an appropriate Taylor’s series.


1
Expert's answer
2022-02-15T11:25:02-0500

Question contains mistake. We have to approximate the value of "\sqrt{10}."



Taylor series is given as

"f(a+h)=f(a)+h.f'(a)+{h^2\over 2!}f''(a)+{h^3\over 3!}f'''(a)+..."

Let's Rewrite "10" in the following form:

"10=9+1"

Now, let's have a function that will enable us use Taylor series method

"f(10)=f(9+1)"

"\implies" "a=9" and "h=1"

Since we are asked to find the square root

"\implies f(a)=a^{1\over 2} \implies f(a+1)=\sqrt {10}"

"\implies f(a)=a^{1\over 2} \implies f(a+1)=\sqrt {10}"


Substituting the value of "a" we have

"f(a)=9^{1\over 2} =3"

"f'(a)={1\over 2}a^{{1\over 2}-1}={1\over 2}a^{-{1\over 2}}={{1\over 2}\over a^{1\over2}}={1\over 2a^{1\over 2}}"

Substituting the value of "a" we have

"f'(a)={1\over 2(9)^{1\over 2}}=0.166..."

"f''(a)=(-{1\over 2})({1\over 2})a^{-{1\over 2}-1}=-{1\over 4}a^{-{3\over 2}}={-{1\over 4}\over a^{3\over 2}}=-{1\over 4a^{3\over 2}}"

Substituting the value of "a" we have

"f''(a)=-{1\over 4(9)^{3\over 2}}=0.00925"

"f'''(a)=({-{3\over 2}})({-{1\over 4}})a^{-{3\over 2}-1}={3\over 8}a^{-{5\over 2}}={{3\over 8}\over a^{5\over 2}}={3\over 8a^{5\over 2}}"

Substituting the value of "a" we have

"f'''(a)={3\over 8(9)^{5\over 2}}=0.000514"

From Taylor series method, we have

"\sqrt {10}=f(a+1)=f(a)+1.f'(a)+{1^2\over 2!}f''(a)+{1^3\over 3!}f'''(a)+..."

"\sqrt {10}=f(a+1)=f(a)+1.f'(a)+{1^2\over 2!}f''(a)+{1^3\over 3!}f'''(a)+..."

But

"f(a)=3" , "f'(a)=0.166.." , "f''(a)=0.00925" , "f'''(a)=0.000514"

"\implies \sqrt {10}= f(a+1)=5+0.1+{0.00925\over 2}+{0.000514\over 3!}=5.10471"

"\implies \sqrt {10}= f(a+1)=5+0.1+{0.00925\over 2}+{0.000514\over 3!}=5.10471"


"\therefore" "\sqrt{10}=5.10471"

"\sqrt{10}=5.10471"

"\sqrt{10}=5.1047"

"\sqrt{10}=5.1047" correct to four decimal places